OpenStudy (anonymous):

rewrite 30/50 as a percent

OpenStudy (anonymous):

60%

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

30/50 = 3/5. 5 x 20 = 100. 3 x 20 = 60.

geerky42 (geerky42):

\(\dfrac{30}{50}\rightarrow\dfrac{60}{100} = 60%\) Of course she is right.

OpenStudy (anonymous):

it was worng

geerky42 (geerky42):

60%*

OpenStudy (anonymous):

30/50= .60 = 60%

OpenStudy (anonymous):

I ment 37/50 my bad

OpenStudy (anonymous):

74%

OpenStudy (anonymous):

what is 4/5 as a percent

geerky42 (geerky42):

Just get denominator to 100, then what in numerator is your answer

OpenStudy (anonymous):

80%

OpenStudy (anonymous):

what is 8% as a decimal

OpenStudy (anonymous):

.08

OpenStudy (anonymous):

85% as a decimal

OpenStudy (anonymous):

.85

OpenStudy (anonymous):

18% as a decimal

OpenStudy (anonymous):

.18

OpenStudy (anonymous):

do you see the pattern ?

OpenStudy (anonymous):

you can do this on your own

OpenStudy (anonymous):

is 6% .6 as a decimal

OpenStudy (anonymous):

.06

OpenStudy (anonymous):

.6 is 60%

OpenStudy (anonymous):

why do you add the 0

OpenStudy (anonymous):

because you are moving the decimal to the right to make a percent

OpenStudy (anonymous):

so 13/20 would be 13%

OpenStudy (anonymous):

type in your calculator 13 divided by 20

OpenStudy (anonymous):

I got 0.65

OpenStudy (anonymous):

now move the decimal over 2 times to the right

OpenStudy (anonymous):

you get 65%

OpenStudy (anonymous):

that's 65.

OpenStudy (anonymous):

also known as multiplying it by 100

OpenStudy (anonymous):

so 43% would be 0.43

OpenStudy (anonymous):

yup !

OpenStudy (anonymous):

and 3/5 would be 6%

OpenStudy (anonymous):

60% if it was .06 then yes it would be 6% Remember you are moving the decimal 2 times

OpenStudy (anonymous):

thank you for your help I really suck at math and I need to know this to graduate

OpenStudy (anonymous):

yup np anytime

OpenStudy (anonymous):

what is 2% of 2.15

OpenStudy (anonymous):

and can you tell me how you got the answer I don't want yall doing my homework for me

OpenStudy (anonymous):

2% multiplied by 2.15

OpenStudy (anonymous):

now you tell me what is 2% as a decimal ?

OpenStudy (anonymous):

.02

OpenStudy (anonymous):

or .2

OpenStudy (anonymous):

first one is correct second one is 20%

OpenStudy (anonymous):

thank you I got it now

OpenStudy (anonymous):

so in your calculator do .02 times 2.15

OpenStudy (anonymous):

ok good so you get it now ?

OpenStudy (anonymous):

yes its easier doing it yourself then asking you to do my work for me

OpenStudy (anonymous):

umm ok